A snowboarder on a slope starts from rest and reaches a speed of 1.7 m/s after 6.5 s. 


 (a) What is the magnitude of the snowboarder's average acceleration?  

 (b) How far does the snowboarder travel in this time? 
Answer #1

Solution,

A)

Acceleration, a = (v - u)/t = (1.7 - 0)/6.5 = 0.26 m/s^2

B)

Distance, s = (v^2 - u^2)/2a = (1.7^2)/(2 x 0.26) = 5.56 m

Comment in case any doubt please rate my answer ....